OpenStudy (anonymous):

What is the equation of a line which includes points (-5, 5) and (2, 5)?

OpenStudy (aravindg):

use 2 point form :)

OpenStudy (sw050399):

Well, you could start with point slope form. y - y1 = m(x - x2) where the point is (x1, y1) and m is the slope.

OpenStudy (aravindg):

where (x1,y1) and (x2,y2) are the points

OpenStudy (anonymous):

Use two point slope form: \[\color{green}{\frac{y - y_1}{y_2 - y_1} = \frac{x - x_1}{x_2 - x_1}}\] \[x_1 = -5, \quad x_2 = 2, \quad y_1 = 5 \quad y_2 = 5\]

OpenStudy (anonymous):

From my close look, these are the two points on the same line.. So, y = 5 should be the line..

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

You are welcome dear..